Please help me to solve.
e^(4-5x)=988
Thank you!

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

e^(4-5x)=988
-----
Take the natural log of both sides to get:
(4-5x) = ln(988)
4-5x = 6.8957
-5x = 2.8957
x = -0.5791
===============
Cheers,
Stan H.
===============

Answer by solver91311(24713)





Take the natural log of both sides:



Use:





Then use:





is just a constant, so solve like any other linear equation in

John